(7%) Problem 5: Suppose you have the 7.75 F capacitor of a heart defibrillator at a potential difference of 17.5 x 104 V. Randomized Variables V-17.5 x 104 v 50% Part (a) What is the energy stored in it in J? U 118671.88 U-118700 Correct! | 50% Part (b) Find the amount of stored charge in mC. Grade Summary Deductions 0% Potential Q = 13.561 100%(7%) Problem 15: select part Charges move in an electric field due to the Coulomb force Some unknown charge moves in a direction perpendicular to the direction of the electric field lines. We can conclude that Grade Summary Deductions Potential Both its electric potential and clectric potential energy stay constant 0% 100% It moves from a high potential to a low potential, and its electric energy decreases It moves from a low potential to a high potential, and its electric energy increases It moves from a high potential to a low potential, and its electric energy increases It moves from a low potential to a high potential, and its electric energy decreases Submissions Attempts remaining: 1 (0% per attempt) detailed view Submit Hint Feedback I give up! 0% 0% Hints: 1 % deduction per hint. Hints remaining: 2 Feedback: 2 for a 0% deduction
Answer #1

(2) Energy = 4 Cr² = 1 * 7.75 X 106x (17.5 x 104 )? = 118672 J Answer Q = cv = 7.75 x 10-6 x 17.58 107 = 1356.25 mc Answer